Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A jar contains 16 red balls and 8 white balls. If 3 balls ar

Expert replies
by Max@Math Revolution » Wed Aug 29, 2018 11:47 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

[Math Revolution GMAT math practice question]

A jar contains 16 red balls and 8 white balls. If 3 balls are selected at random from the jar, what is the approximate probability that all balls selected are white?

A. 0.02
B. 0.03
C. 0.04
D. 0.05
E. 0.06
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Thu Aug 30, 2018 2:25 am
Max@Math Revolution wrote:[Math Revolution GMAT math practice question]

A jar contains 16 red balls and 8 white balls. If 3 balls are selected at random from the jar, what is the approximate probability that all balls selected are white?

A. 0.02
B. 0.03
C. 0.04
D. 0.05
E. 0.06
P(1st marble is white) = 8/24. (Of the 24 marbles, 8 are white.)
P(2nd marble is white) = 7/23. (Of the 23 remaining marbles, 7 are white.)
P(3rd marble is white) = 6/22. (Of the 22 remaining marbles, 6 are white.)
To combine these probabilities, we multiply:
8/24 * 7/23 * 6/22 = 1/3 * 7/23 * 3/11 = 7/(23*11) = 7/253 ≈ 7/250 = 28/1000 = 0.28 ≈ 0.03.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

Max@Math Revolution wrote:[Math Revolution GMAT math practice question]

A jar contains 16 red balls and 8 white balls. If 3 balls are selected at random from the jar, what is the approximate probability that all balls selected are white?

A. 0.02
B. 0.03
C. 0.04
D. 0.05
E. 0.06
\[{\text{Jar}}\,\,\left\{ \begin{gathered}
16\,r \hfill \\
8\,w \hfill \\
\end{gathered} \right.\,\,\,\,\, \Rightarrow \,\,\,3\,\,{\text{simultaneous}}\,\,{\text{extractions}}\]
\[{\text{? = P}}\left( {{\text{all}}\,\,w} \right)\]

\[{\text{total}} = C\left( {16 + 8,3} \right) = C\left( {24,3} \right)\,\,\,\,{\text{equiprobable}}\]
\[{\text{favorable}} = C\left( {8,3} \right)\]

\[? = \frac{{C\left( {8,3} \right)}}{{C\left( {24,3} \right)}} = \cdots = \frac{7}{{253}}\]
\[{\left( ? \right)^{ - 1}} = \frac{{210 + 42 + 1}}{7} = \boxed{36\frac{1}{7}}\]
\[{\left( A \right)^{ - 1}} = {\left( {\frac{2}{{100}}} \right)^{ - 1}} = 50\]
\[{\left( B \right)^{ - 1}} = {\left( {\frac{3}{{100}}} \right)^{ - 1}} = \frac{{99 + 1}}{3} = \boxed{33\frac{1}{3}}\]
From the fact that 33 1/3 is less than 36 1/7 , we are sure the other alternative choices (all of them are in increasing order) are not closer to our FOCUS.
(In other words, the reciprocals of (C), (D) and (E) are certainly less than 33 1/3 , hence the closest to 36 1/7 is 33 1/3.)

The above follows the notations and rationale taught in the GMATH method.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

by Brent@GMATPrepNow » Thu Aug 30, 2018 6:46 am
Max@Math Revolution wrote:[Math Revolution GMAT math practice question]

A jar contains 16 red balls and 8 white balls. If 3 balls are selected at random from the jar, what is the approximate probability that all balls selected are white?

A. 0.02
B. 0.03
C. 0.04
D. 0.05
E. 0.06
There are 24 balls in the jar (before any are removed)

P(all 3 balls are white) = P(1st ball is white AND 2nd ball is white AND 3rd ball is white)
= P(1st ball is white) x P(2nd ball is white) x P(3rd ball is white)
= 8/24 x 7/23 x 6/22
= 1/3 x 7/23 x 3/11
= 7/253

ASIDE: rather than perform long division, we can take 7/253 and create an equivalent fraction by multiplying top and bottom by 4 to get...
7/253 = (7)(4)/(253)(4)
≈ 28/1000
≈ 0.028
≈0.03

Answer: B
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

Brent@GMATPrepNow wrote: ASIDE: rather than perform long division, we can take 7/253 and create an equivalent fraction by multiplying top and bottom by 4 to get...
7/253 = (7)(4)/(253)(4)
≈ 28/1000
≈ 0.028
≈0.03
Long division is something that must be avoided. That´s why GMATH´s approach deals with what we call "breaking numbers".

It is not comfortable to do (say) 253/7 by long division - and we did not - this can be done WITHOUT APPROXIMATIONS easily, as we did and the way we repeat below:

253 = 210 + 42 + 1 , this choice is "smart" because 210 and 42 are not only divisible by 7 (obviously), but also because their quotients are trivial (30 and 6)...

Therefore 253/7 = 210/7 + 42/7 + 1/7 = 36 + 1/7 , without FULL PRECISION!

One last detail: when we approximate (say) 7/253 by 7/250 , we must be cautious because approximations in denominators may have "small errors" that propagate according to the "magnitude" (absolute value) of the numerator... that´s why we believe the approach we have taken is safer, although in this case the approximation 7/253 by 7/250 was sufficiently good for the alternative choices.

The above follows the notations and rationale taught in the GMATH method.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

edit

by Max@Math Revolution » Sun Sep 02, 2018 5:35 pm
=>

There are 24C3 ways of choosing 3 balls from the 24 in the jar, and 8C3 ways of choosing 3 balls from the 8 white balls. Therefore, the probability of choosing 3 white balls from the jar is:
8C3 / 24C3 = { (8*7*6) / (1*2*3) } / { (24*23*22) / (1*2*3) } = (8*7*6) / (24*23*22) = 7 / ( 23* 11 ) ≈ 1/ (3*11) ≈ 0.03

Therefore, the answer is B.
Answer: B
Join the discussion